1. Job cost sheets show the following information:

Job January February March Completed Sold
AA2 $2,400     $1,000     February Not sold
AA4 4,836     January February
AA5 3,260     February March
AA3 3,409     $2,320 April Not sold
Total $7,236     $7,669     $2,320

What are the balances in the work in process inventory, finished goods inventory, and cost of goods sold for January, February, and March?

Work in
Process
Finished
Goods
COGS
January $2400 $4836 0
February 3409 ??? $4836
March 5729 ??? 3260

***Need answers for ???

2. Pocono Cement Forms expects $1,050,000 in overhead during the next year. It does not know whether it should apply overhead on the basis of its anticipated direct labor hours of 70,000 or its expected machine hours of 35,000.

Determine the product cost under each predetermined allocation rate if the last job incurred $1,530 in direct material cost, 84 direct labor hours, and 70 machine hours. Wages are paid at $16 per hour.

Labor Hours Machine hours
Cost of the job $4134 $

Need answer for Machine Hours
Answer #1

1.
February - Finished Goods = $2400+1000+3260 = $6660
March - Finished Goods = $2400+1000 = $3400

2.
Using machine hours
Overhead rate = $1050000 / 35000 = $30 per machine hour
Cost of the job = $1530 + 84 x 16 + 70 x 30 = $4974

If you have any query, kindly comment with your query and please mark thumbs up.